# ProtTeX: Structure-In-Context Reasoning and Editing of Proteins with Large Language Models

### v2

@@ 1.   Introduction
Large language models (LLMs) exhibit scalability, emergent abilities, and generality, enabling them to transcend the limitations of single-task models and facilitate cross-domain knowledge transfer11, 12. They are fundamentally reshaping the paradigm of scientific research and serve as a highly suitable unified framework for multitask learning. In the realm of small molecules, advanced LLMs have demonstrated remarkable capabilities in predicting molecular properties, understanding functional characteristics, and designing novel molecules. For instance, Chen et al.13 developed MatterGPT, a generative transformer model that employs the SLICES (Simplified Line-Input Crystal-Encoding System) 14 representation to achieve on-demand inverse design of solid-state materials with single and multiple targeted properties. The success of LLMs in this domain can be attributed to the use of domain-specific tokenizers, such as SMILES (Simplified Molecular Input Line Entry System) 15, SELFIES (SELF-referencIng Embedded Strings) 16, and SLICES, which effectively encode molecular representations for learning and inference. However, in the field of proteins, tokenization has been largely limited to the use of one-letter amino acid representations, which prevents LLMs from effectively handling structure-related tasks. For example, Wang et al. 17 introduced InstructProtein, enabling the LLM with bidirectional generation capabilities between human language and protein sequences. TourSynbio18 fine-tunes the InternLM2-7B19 utilizing the ProteinLMBench 20 dataset to augment its capability in comprehending protein sequences. Tokenizing protein sequences using one-letter abbreviations leads to ambiguity with textual characters and often results in mismatches between amino acid length and tokenized sequence length, which can obscure the semantic representation of sequence elements. Furthermore, sequence-only representations are insufficient for fully understanding proteins, as functional inference often requires structural information. To tackle this challenge, many approaches employ compositional modality-specific protein encoders 21, 22, 23, 24, 25. These models face optimization challenges and are limited to text generation, struggling with multimodal generation and reasoning. Moreover, Several purported LLM-based models 24, 17 still rely on classification-based protein function prediction, restricting prompt flexibility and model adaptability. In biological research, function classification is inherently open-ended due to the continuous discovery of novel proteins and the evolving nature of taxonomies. For instance, in the Gene Ontology database 26, the number of molecular function terms for humans increased by 26% in 2018 compared to 201627. Given LLM architectures’ strengths in generative modeling, we emphasize generative capabilities over rigid classification paradigms. Overly constrained prompts and classification heads would limit the model’s ability to handle novel or ambiguous functions, undermining the goal of developing a robust and adaptive protein chat system. Current models also exhibit a significant limitation in their inability to perform reasoning, e.g. the Chain-of-Thought (CoT) approach 28, which introduces a sequence of intermediate reasoning steps within in-context examples, facilitating complex reasoning processes in LLMs. In protein science, understanding or annotating proteins often necessitates the derivation and inference of information from sequences, structures, and other relevant data. To enable LLMs to perform reasoning and deduction on protein structures, the introduction of an effective structure-based tokenizer is essential. However, the integration of such a tokenizer into LLMs and the exploration of reasoning abilities remain unexplored.
@@
Inspired by the multi-modal LLMs Emu3 29 and Chamelon 30, we present ProtTeX, a dedicated framework for tokenizing both the 1D sequence and 3D structure of proteins for LLMs, akin to SMILES. Just as TeX provides precise control over document layout and formatting, ProtTeX empowers researchers for protein data formatting and editing, which enables LLMs to reason and generate sequences that arbitrarily interleave textual and protein modality. ProtTeX can employ the advanced Reasoning paradigm, CoT reasoning, to enhance LLMs’ capability in protein-related tasks. This approach allows structural or textual information to serve as a key logical component in the reasoning process, making protein problem solving more transparent, logical, interpretable, and controllable. We have constructed a unified foundational framework that leverages a single model and loss function based solely on the next-token prediction (NTP) strategy, enabling seamless adaptation to diverse protein-related downstream tasks. In comparison to the state-of-the-art domain expert LLMs BiomedGPT 23 and Llama2-molinst-protein-7B 31, our approach demonstrates a twofold enhancement in accuracy for protein function prediction. Our findings reveal that the CoT approach and sampling strategies from the field of LLM can be effortlessly adapted to the protein domain, enabling high-accuracy protein structure prediction and conformational sampling. Additionally, the integration of arbitrary textual inputs allows for controllable protein generation based on human-defined prompts. For the first time, we demonstrate that decoder-only LLMs are capable of understanding, predicting, and designing proteins.
@@ 2.1.   Tokenizing All-Atom Protein Structures
Then, we use the original amino acid sequence as our protein side-chain tokenizer. Instead of directly using abbreviation letters or other compositional encoders, we simply add 20 new special tokens to LLMs to represent the protein sequence, similar to Emu329. We reinitialize 512+20=532 new tokens using the same methodology as that applied to the original textual tokens.
@@ 2.3.1 Dataset
The protein-related QA pairs were curated from Mol-Instruction31 and ProteinLMBench20, with all protein accessions sourced from UniProt 42. By mapping the corresponding accessions of specific QA pairs from the training, validation, and test sets, we constructed three distinct datasets: Protein Function Understanding Dataset (PFUD), Protein Structure Analysis Dataset (PSAD), and Protein Design Dataset (PDD). Among these, PFUD and PDD were derived and modified from the Mol-Instruction dataset, while PSAD was derived from a subset of ProteinLMBench. If an accession does not exist in our database, the corresponding questions were dropped. The remaining proteins without corresponding QA pairs formed the Protein Structure Prediction Dataset (PSPD). The total dataset for training our main model is composed of the aforementioned four components—PFUD, PSAD, PDD, and PSPD—and is randomly shuffled at the beginning of each epoch to ensure robust training and prevent any potential bias introduced by the order of the data. The token counts for each subset are presented in Table 1. For more details on our dataset, please see the supplementary section A.2.
@@ 2.3.2 Training
Instead of training a completely new model from scratch, we opt for continuous pre-training 43 and supervised fine-tuning 44 of existing general LLMs. We assign equal weight to the tokens of the protein modality and natural language, given the critical importance of protein information. For further details on training, please refer to the supplementary section A.3.
@@ 2.3.3 Inference and sampling
The nucleus sampling strategy 45 is used for the multi-conformation generation and protein design tasks.
@@ 2.3.4 Baseline Setups
  • Llama3-Instruct Meta-Llama-3-8B-Instruct 46 is an advanced language model developed by Meta, featuring 8 billion parameters and optimized for instructional fine-tuning to enhance conversational performance. Despite its relatively smaller parameter size compared to other models, the Meta-Llama-3-8B-Instruct has demonstrated superior performance, even surpassing the 70-billion-parameter Llama-2 model 47 in various benchmarks.
  • BioMedGPT-LM-10B 23 BioMedGPT is a domain-specific LLM fine-tuned on a large selection of biological scientific corpora, including protein-related questions. The model encodes the protein sequence with ESM-3B and uses BioMedGPT-LM-7B as the decoder to generate responses.
  • Llama2-molinst-protein-7B This model was developed by the research group of the Mol-Instructions 31 dataset. They performed full-parameter fine-tuning of Llama2-7B 47 using the protein-oriented dataset in Mol-Instructions. For our inference, identical parameters to those specified in the official scripts were employed.
@@
  • ProtT3-FT ProtT3 25 empowers an LLM to understand protein sequences by incorporating ESM-3B 48 as its protein understanding module, enabling effective protein-to-text generation. Although it has undergone large-scale pretraining and fine-tuning on protein-text retrieval and generation, it imposes rigid constraints on its question templates, lacking support for diverse prompt variations and questions, which represents a significant limitation and undermines its claim as a genuine multimodal large model. To address this critical issue, we fine-tuned the model using our PFUD dataset and selected the best model in benchmark evaluations, which resulted in substantial performance improvements and enhanced model capabilities.
  • $\text{ProtTeX}_{\text{Llama3}}$ Our main proposed model, which is supervised fine-tuned on Meta-Llama-3-8B-Base using the ProtTeX tokenizer with our total dataset, which incorporates training data from different downstream tasks. The textual protein sequence information in PFUD dataset is replaced by our tokenized protein sequence and structure. We also conducted ablation experiments under different training conditions, including various training strategies such as Low-Rank Adaptation (LoRA) 49, different training sets, and model scales. For more details on the ablation experiments, please refer to Table 4, Table 3, and Section A.4.
@@ 2.3.5 Metrics
Traditional multi-class classification models for protein function prediction rely on the CAFA 50 evaluation metrics. These metrics require the model to output scores for each classification head to compute the Fmax. However, these metrics are not applicable to large language generation models, as we do not employ a classification-based approach and thus do not output scores for individual classes. Consequently, we have developed a novel evaluation framework tailored to these models. We evaluate the model’s output based on two key aspects: fluency and domain-specific accuracy. First, we employ two classical metrics, BLEU 51 and ROUGE 52, which are widely used for evaluating machine translation quality. These metrics measure the overlap between machine-generated text and reference translations by comparing n-grams.
@@
Due to the diversity of labels in the original dataset and generated output, manually extracting keywords is highly challenging. To address this, we use DeepSeek-V3 53 to automatically extract keywords and perform exact matching using in-context learning, thereby computing the EMJI. This approach ensures a more efficient and consistent evaluation process while maintaining high precision in keyword matching.
@@ 3.1.   ProtTeX Enables Structure-In-Context Protein Understanding
For our comparative analysis, we curate single-turn dialogue samples from the PFUD test set, resulting in a collection of 5,836 distinct protein queries. These single-turn dialogues encompass six distinct domains of inquiry pertaining to proteins: molecular function, subcellular location, biological process, domains or motifs, overview of features, and multi-attribute. Specifically, the "overview of features" domain prompts the model to provide a concise summary of the most critical functional aspects of the protein, which may encompass elements of both molecular function and biological process. Within the "multi-attribute" domain, our prompts are designed to query combinations of two or three attributes from the preceding five domains, requiring the model to address them collectively in its response. We select three open-source models for our benchmark: BioMedGPT-LM-10B (BioMedGPT) 23, a domain-specific expert model finetuned on protein-related questions. Llama2-molinst-protein-7B, a state-of-the-art protein understanding LLM fine-tuned on Mol-Instructions. Llama3-Instruct (8B) 46, a general-purpose instruction tuned LLM. To compare our model with other compositional models, we also fine-tuned a state-of-the-art compositional protein understanding model, ProtT3 25, as the original ProtT3 constrains the prompts and could not be adapted to other questions. For more details on the models, please refer to the baseline sections 2.3.4.
@@
As shown in Table 2, our analysis demonstrates that the multi-task fine-tuned model $\text{ProtTeX}_{\text{Llama3}}$ achieves optimal performance, excelling in both linguistic fluency and accuracy in addressing domain-specific professional protein queries, highlighting the effectiveness of our ProtTeX framework and the existence of overlapping subspaces among different tasks which facilitate mutual enhancement during the training process. Furthermore, we conduct a comparative analysis by fine-tuning models exclusively on the PFUD dataset to evaluate their respective performance in the ablation study A.4. The experimental results in Table 2 and Table 4 demonstrate that incorporating the novel ProtTeX tokenizer significantly outperforms the the fine-tuned models that use one-letter amino acid abbreviations (Llama2-molinst-protein-7B,Llama3-AAseq-FT) or compositional protein encoders (ProtT3-FT, BiomedGPT). These results highlights the limitations of current models in functional understanding when relying solely on textual sequence alphabets or other compositional approaches, further substantiating the beneficial impact of incorporating structural tokens in enhancing protein function comprehension. On our PFUD test set, where all QA pairs are derived exclusively from Swiss-Prot proteins, the ROUGE-L score of the Llama2-molinst-protein-7B 31 model closely matches its officially reported metrics. However, we find that ROUGE-L is not a robust indicator for assessing the precision of key terms, as demonstrated by the substantially lower exact match performance. Our ablation studies presented in Table 4 demonstrate that pre-training without explicit functional information significantly enhances functional comprehension. These results collectively suggest a symbiotic relationship between multimodal understanding and generation. Specifically, enhancing comprehension data improves generative task performance, and conversely, expanding generative training data strengthens interpretative capabilities. This phenomenon aligns with recent observations in computer vision research 54. Additionally, we evaluated domain-specific response performance, as illustrated in Figure 2. Although the fine-tuned ProtT3 model achieves comparable performance with our proposed model in motif recognition and multi-attribute tasks, ProtTeX consistently outperforms competing approaches across various domains. Notably, the BiomedGPT model exhibits suboptimal performance throughout our experiments. Despite extensive efforts to optimize the prompt formulation by adopting BiomedGPT’s preferred structure, the model demonstrates substantial limitations in addressing a wide range of protein-related queries. This suggests that BiomedGPT’s training may be insufficient for comprehensive protein-related tasks. Moreover, approximately 2% of BiomedGPT’s responses yielded unknown answers, indicating that the external encoder integration strategy may cause substantial representational shifts in response to sequence similarity variations, potentially increasing perplexity and reducing overall performance. These observations underscore the effectiveness of our early fusion training strategy 55, 56, 57, which integrates interleaved textual and protein modalities into a unified representation, enabling natural and intrinsic connections between different modalities.

### v3

@@ 1.   Introduction
Large language models (LLMs) exhibit scalability, emergent abilities, and generality, enabling them to transcend the limitations of single-task models and facilitate cross-domain knowledge transfer11, 12. They are fundamentally reshaping the paradigm of scientific research and serve as a highly suitable unified framework for multitask learning. In the realm of small molecules, advanced LLMs have demonstrated remarkable capabilities in predicting molecular properties, understanding functional characteristics, and designing novel molecules. For instance, Chen et al.13 developed MatterGPT, a generative transformer model that employs the SLICES (Simplified Line-Input Crystal-Encoding System) 14 representation to achieve on-demand inverse design of solid-state materials with single and multiple targeted properties. The success of LLMs in this domain can be attributed to the use of domain-specific tokenizers, such as SMILES (Simplified Molecular Input Line Entry System) 15, SELFIES (SELF-referencIng Embedded Strings) 16, and SLICES, which effectively encode molecular representations for learning and inference. In the field of proteins, tokenization has been largely limited to the use of one-letter amino acid representations 17, 18, 19. For example, Llama2-molinst-protein-7B 19 is fine-tuned from the Llama2-7B model using the protein-oriented dataset from Mol-instructions 19, enabling diverse protein function understanding. However, tokenizing protein sequences using one-letter abbreviations leads to ambiguity with textual characters and often results in mismatches between amino acid length and tokenized sequence length, which can obscure the semantic representation of sequence elements. Furthermore, sequence-only representations are insufficient for fully understanding proteins, as functional inference often requires structural information. To tackle this challenge, many approaches employ compositional modality-specific protein encoders 20, 21, 22, 23, 24. For example, BioMedGPT-LM-10B 22 employs ESM2-3B 25 as its encoder and is fine-tuned from the Llama2-7B 26 model using millions of protein-text question-answering pairs, enabling the generation of natural language descriptions for proteins. However, these models face optimization challenges and are limited to text generation, struggling with multimodal generation and reasoning. Moreover, Several purported LLM-based models 23, 17 still rely on classification-based protein function prediction, restricting prompt flexibility and model adaptability. In biological research, function classification is inherently open-ended due to the continuous discovery of novel proteins and the evolving nature of taxonomies. For instance, in the Gene Ontology database 27, the number of molecular function terms for humans increased by 26% in 2018 compared to 201628. Given LLM architectures’ strengths in generative modeling, we emphasize generative capabilities over rigid classification paradigms. Overly constrained prompts and classification heads would limit the model’s ability to handle novel or ambiguous functions, undermining the goal of developing a robust and adaptive protein chat system. Current models also exhibit a significant limitation in their inability to perform reasoning, e.g. the Chain-of-Thought (CoT) approach 29, which introduces a sequence of intermediate reasoning steps within in-context examples, facilitating complex reasoning processes in LLMs. In protein science, understanding or annotating proteins often necessitates the derivation and inference of information from sequences, structures, and other relevant data. To enable LLMs to perform reasoning and deduction on protein structures, the introduction of an effective structure-based tokenizer is essential. However, the integration of such a tokenizer into LLMs and the exploration of reasoning abilities remain unexplored.
@@
Inspired by the multi-modal LLMs Emu3 30 and Chamelon 31, we present ProtTeX, a dedicated framework for tokenizing both the 1D sequence and 3D structure of proteins for LLMs, akin to SMILES. Just as TeX provides precise control over document layout and formatting, ProtTeX empowers researchers for protein data formatting and editing, which enables LLMs to reason and generate sequences that arbitrarily interleave textual and protein modality. ProtTeX can employ the advanced Reasoning paradigm, CoT reasoning, to enhance LLMs’ capability in protein-related tasks. This approach allows structural or textual information to serve as a key logical component in the reasoning process, making protein problem solving more transparent, logical, interpretable, and controllable. We have constructed a unified foundational framework that leverages a single model and loss function based solely on the next-token prediction (NTP) strategy, enabling seamless adaptation to diverse protein-related downstream tasks. In comparison to the state-of-the-art domain expert LLMs BiomedGPT 22 and Llama2-molinst-protein-7B 19, our approach demonstrates a twofold enhancement in accuracy for protein function prediction. Our findings reveal that the CoT approach and sampling strategies from the field of LLM can be effortlessly adapted to the protein domain, enabling high-accuracy protein structure prediction and conformational sampling. Additionally, the integration of arbitrary textual inputs allows for controllable protein generation based on human-defined prompts. For the first time, we demonstrate that decoder-only LLMs are capable of understanding, predicting, and designing proteins.
@@ 2.1.   Tokenizing All-Atom Protein Structures
Then, we use the original amino acid sequence as our protein side-chain tokenizer. Instead of directly using abbreviation letters or other compositional encoders, we simply add 20 new special tokens to LLMs to represent the protein sequence, similar to Emu330. We reinitialize 512+20=532 new tokens using the same methodology as that applied to the original textual tokens.
@@ 2.3.1 Dataset
The protein-related QA pairs were curated from Mol-Instruction19 and ProteinLMBench42, with all protein accessions sourced from UniProt 43. By mapping the corresponding accessions of specific QA pairs from the training, validation, and test sets, we constructed three distinct datasets: Protein Function Understanding Dataset (PFUD), Protein Structure Analysis Dataset (PSAD), and Protein Design Dataset (PDD). Among these, PFUD and PDD were derived and modified from the Mol-Instruction dataset, while PSAD was derived from a subset of ProteinLMBench. If an accession does not exist in our database, the corresponding questions were dropped. The remaining proteins without corresponding QA pairs formed the Protein Structure Prediction Dataset (PSPD). The total dataset for training our main model is composed of the aforementioned four components—PFUD, PSAD, PDD, and PSPD—and is randomly shuffled at the beginning of each epoch to ensure robust training and prevent any potential bias introduced by the order of the data. The token counts for each subset are presented in Table 1. For more details on our dataset, please see the supplementary section A.2.
@@ 2.3.2 Training
Instead of training a completely new model from scratch, we opt for continuous pre-training 44 and supervised fine-tuning 45 of existing general LLMs. We assign equal weight to the tokens of the protein modality and natural language, given the critical importance of protein information. For further details on training, please refer to the supplementary section A.3.
@@ 2.3.3 Inference and sampling
The nucleus sampling strategy 46 is used for the multi-conformation generation and protein design tasks.
@@ 2.3.4 Baseline Setups
  • Llama3-Instruct Meta-Llama-3-8B-Instruct 47 is an advanced language model developed by Meta, featuring 8 billion parameters and optimized for instructional fine-tuning to enhance conversational performance. Despite its relatively smaller parameter size compared to other models, the Meta-Llama-3-8B-Instruct has demonstrated superior performance, even surpassing the 70-billion-parameter Llama-2 model 26 in various benchmarks.
  • BioMedGPT-LM-10B 22 BioMedGPT is a domain-specific LLM fine-tuned on a large selection of biological scientific corpora, including protein-related questions. The model encodes the protein sequence with ESM-3B and uses BioMedGPT-LM-7B as the decoder to generate responses.
  • Llama2-molinst-protein-7B This model was developed by the research group of the Mol-Instructions 19 dataset. They performed full-parameter fine-tuning of Llama2-7B 26 using the protein-oriented dataset in Mol-Instructions. For our inference, identical parameters to those specified in the official scripts were employed.
@@
  • ProtT3-FT ProtT3 24 empowers an LLM to understand protein sequences by incorporating ESM-3B 25 as its protein understanding module, enabling effective protein-to-text generation. Although it has undergone large-scale pretraining and fine-tuning on protein-text retrieval and generation, it imposes rigid constraints on its question templates, lacking support for diverse prompt variations and questions, which represents a significant limitation and undermines its claim as a genuine multimodal large model. To address this critical issue, we fine-tuned the model using our PFUD dataset and selected the best model in benchmark evaluations, which resulted in substantial performance improvements and enhanced model capabilities.
  • $\text{ProtTeX}_{\text{Llama3}}$ Our main proposed model, which is supervised fine-tuned on Meta-Llama-3-8B-Base using the ProtTeX tokenizer with our total dataset, which incorporates training data from different downstream tasks. The textual protein sequence information in PFUD dataset is replaced by our tokenized protein sequence and structure. We also conducted ablation experiments under different training conditions, including various training strategies such as Low-Rank Adaptation (LoRA) 48, different training sets, and model scales. For more details on the ablation experiments, please refer to Table 4, Table 3, and Section A.4.
@@ 2.3.5 Metrics
Traditional multi-class classification models for protein function prediction rely on the CAFA 49 evaluation metrics. These metrics require the model to output scores for each classification head to compute the Fmax. However, these metrics are not applicable to large language generation models, as we do not employ a classification-based approach and thus do not output scores for individual classes. Consequently, we have developed a novel evaluation framework tailored to these models. We evaluate the model’s output based on two key aspects: fluency and domain-specific accuracy. First, we employ two classical metrics, BLEU 50 and ROUGE 51, which are widely used for evaluating machine translation quality. These metrics measure the overlap between machine-generated text and reference translations by comparing n-grams.
@@
Due to the diversity of labels in the original dataset and generated output, manually extracting keywords is highly challenging. To address this, we use DeepSeek-V3 52 to automatically extract keywords and perform exact matching using in-context learning, thereby computing the EMJI. This approach ensures a more efficient and consistent evaluation process while maintaining high precision in keyword matching.
@@ 3.1.   ProtTeX Enables Structure-In-Context Protein Understanding
For our comparative analysis, we curate single-turn dialogue samples from the PFUD test set, resulting in a collection of 5,836 distinct protein queries. These single-turn dialogues encompass six distinct domains of inquiry pertaining to proteins: molecular function, subcellular location, biological process, domains or motifs, overview of features, and multi-attribute. Specifically, the "overview of features" domain prompts the model to provide a concise summary of the most critical functional aspects of the protein, which may encompass elements of both molecular function and biological process. Within the "multi-attribute" domain, our prompts are designed to query combinations of two or three attributes from the preceding five domains, requiring the model to address them collectively in its response. We select three open-source models for our benchmark: BioMedGPT-LM-10B (BioMedGPT) 22, a domain-specific expert model finetuned on protein-related questions. Llama2-molinst-protein-7B, a state-of-the-art protein understanding LLM fine-tuned on Mol-Instructions. Llama3-Instruct (8B) 47, a general-purpose instruction tuned LLM. To compare our model with other compositional models, we also fine-tuned a state-of-the-art compositional protein understanding model, ProtT3 24, as the original ProtT3 constrains the prompts and could not be adapted to other questions. For more details on the models, please refer to the baseline sections 2.3.4.
@@
As shown in Table 2, our analysis demonstrates that the multi-task fine-tuned model $\text{ProtTeX}_{\text{Llama3}}$ achieves optimal performance, excelling in both linguistic fluency and accuracy in addressing domain-specific professional protein queries, highlighting the effectiveness of our ProtTeX framework and the existence of overlapping subspaces among different tasks which facilitate mutual enhancement during the training process. Furthermore, we conduct a comparative analysis by fine-tuning models exclusively on the PFUD dataset to evaluate their respective performance in the ablation study A.4. The experimental results in Table 2 and Table 4 demonstrate that incorporating the novel ProtTeX tokenizer significantly outperforms the the fine-tuned models that use one-letter amino acid abbreviations (Llama2-molinst-protein-7B,Llama3-AAseq-FT) or compositional protein encoders (ProtT3-FT, BiomedGPT). These results highlights the limitations of current models in functional understanding when relying solely on textual sequence alphabets or other compositional approaches, further substantiating the beneficial impact of incorporating structural tokens in enhancing protein function comprehension. On our PFUD test set, where all QA pairs are derived exclusively from Swiss-Prot proteins, the ROUGE-L score of the Llama2-molinst-protein-7B 19 model closely matches its officially reported metrics. However, we find that ROUGE-L is not a robust indicator for assessing the precision of key terms, as demonstrated by the substantially lower exact match performance. Our ablation studies presented in Table 4 demonstrate that pre-training without explicit functional information significantly enhances functional comprehension. These results collectively suggest a symbiotic relationship between multimodal understanding and generation. Specifically, enhancing comprehension data improves generative task performance, and conversely, expanding generative training data strengthens interpretative capabilities. This phenomenon aligns with recent observations in computer vision research 53. Additionally, we evaluated domain-specific response performance, as illustrated in Figure 2. Although the fine-tuned ProtT3 model achieves comparable performance with our proposed model in motif recognition and multi-attribute tasks, ProtTeX consistently outperforms competing approaches across various domains. Notably, the BiomedGPT model exhibits suboptimal performance throughout our experiments. Despite extensive efforts to optimize the prompt formulation by adopting BiomedGPT’s preferred structure, the model demonstrates substantial limitations in addressing a wide range of protein-related queries. This suggests that BiomedGPT’s training may be insufficient for comprehensive protein-related tasks. Moreover, approximately 2% of BiomedGPT’s responses yielded unknown answers, indicating that the external encoder integration strategy may cause substantial representational shifts in response to sequence similarity variations, potentially increasing perplexity and reducing overall performance. These observations underscore the effectiveness of our early fusion training strategy 54, 55, 56, which integrates interleaved textual and protein modalities into a unified representation, enabling natural and intrinsic connections between different modalities.